OpenStudy (anonymous):

What is the equation of the line that passes through the point (1, 6) and has a slope of 3? A) y = 3x + 5 B) y = 3x + 3 C) y = 3x - 17 D) y = 3x - 5

OpenStudy (anonymous):

@Notamathgenius

OpenStudy (notamathgenius):

Well Y=6 so take 1 and multiply them all by 1 and see which one gives 6

OpenStudy (johnweldon1993):

point slope formula y - y1 = m(x - x1) you have a point... (1 , 6) (x , y) and a slope 3 plug them in y - 6 = 3(x - 1) and simplify

OpenStudy (anonymous):

y = mx + b